An indicator is a weak acid and pH range of its colour change is 3.1 to 4.5.if the neutral point of indicator in the center of this pH range,What is ionization constant of the indicator?
1
Expert's answer
2020-07-07T14:37:16-0400

The ionization equilibrium of the indicator:

HA = H+ + A-

ionization constant is: K = [H+]*[A-]/[HA]

at neutral point: [A-] = [HA],

hence, K = [H+]

if the neutral point is in the center of the pH range 3.1-4.5, it is at the pH = (4.5 + 3.1)/2 = 3.8

Thus, ionization constant is K = 10-3.8 = 1.58*10-4